# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ Abstract.
We consider cross-diffusion systems describing evolution of two species $u$ and $v$ moving according to the Darcy’s law with the pressure law $p(s)=\frac{1}{\alpha-1}s^{\alpha-1}$ where $s=u+v$. One of the most challenging questions in the field is the construction of solutions to the problem in the presence of additional advection fields, without imposing any artificial structure (for instance, preventing the species from mixing). Although advection arises naturally in these models, it breaks the symmetry of the system and prevents application of techniques developed in recent years. Here, we provide a new approach to construct solutions in one space dimension that works in a unified manner for all pressure exponents $\alpha\in(0,\infty)$ and for arbitrary initial data (segregated, mixed, or partially mixed). In particular, in the regime $\alpha>1$, this yields the first existence result of its kind, obtained without any structural assumptions. We construct the solutions as a limit of a vanishing viscosity approximation $(u_{\varepsilon},v_{\varepsilon})$. The main challenge is to identify the limit of $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$, where $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$. The key new insight is that possible oscillations of $u_{\varepsilon}$ and $\partial_{x}p(s_{\varepsilon})$ are correlated, simplifying the Young measure analysis in the compensated compactness argument and allowing identification of the limit.
@@ 1. Introduction
Cross-diffusion systems have attracted considerable attention in recent years due to their analytical challenges [21, 33, 17, 6, 30] and their applications in the applied sciences [13, 26, 37]. Here, we consider the following cross-diffusion system
@@
for the evolution of two species on $\mathbb{R}$ having nonnegative densities $u$, $v$, where we denote
@@
$p(s)=\begin{cases}\frac{1}{\alpha-1}s^{\alpha-1}&\mbox{ if }\alpha\neq 1,\\ \log s&\mbox{ if }\alpha=1,\end{cases}\qquad\qquad\mbox{ for }\alpha\in(0,\infty),$
and it models the fact that species avoid overcrowded regions. The case $\alpha>1$ seems to be more relevant from the point of view of applications as the total density $s$ is expected to move with finite speed of propagation by an analogy to the porous medium equation [50, Chap. 14]. On the other hand, for $\alpha\leq 1$ one expects that $s$ is immediately globally supported as for heat equation or fast-diffusion equation, see [31, Theorem 2.4] and [49]. Moreover, the so-called incompressible limit $\alpha\to\infty$ shows that these models are related to the free boundary models of tissue growth for large $\alpha$ [40, 29].
@@
The target of this work is to construct weak solutions without imposing any structural conditions on $V^{i}$, $K^{i,j}$ or without imposing conditions guaranteeing that $u$, $v$ or $u+v$ are strictly positive. In particular, we do not make any assumption on the support of the initial conditions, the pressure law exponent and velocity fields (except mild regularity assumptions which could be probably improved).
@@
with $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$ and study the limit $\varepsilon\to 0$. The difficult terms are nonlinear ones, namely $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$, $v_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$. Without advection (i.e. $\mathcal{V}^{1}=\mathcal{V}^{2}=0$) the system has attracted a large group of researchers and the understanding is now quite complete in arbitrary number of dimensions. In particular, there are several techniques to construct weak solutions: Aronson-Benilan estimates [4] can be used to obtain a uniform bound on $\partial^{2}_{x}p(s_{\varepsilon})$ [30], BV estimates on $\frac{u_{\varepsilon}}{s_{\varepsilon}}$, $\frac{v_{\varepsilon}}{s_{\varepsilon}}$ coming from a transport equation satisfied by these quantities lead to strong compactness of $u_{\varepsilon}$ and $v_{\varepsilon}$ in one dimension (probably this argument was used in [7] for the first time but see [9, 10] as well) while direct energy considerations show that $\{\partial_{x}p(s_{\varepsilon})\}$ is strongly compact in $L^{2}((0,T)\times\mathbb{R}^{d})$ for $\alpha\geq 1$ [19, 43, 24]. In contrast, for a single species the problem with advection is well understood [20].
@@
  • [32] constructs weak solutions to (1.2) with $\alpha>1$ without nonlocal interaction $K^{i,j}=0$. The initial conditions are initially separated and the Authors impose an inequality between the velocities $\partial_{x}V^{1}$, $\partial_{x}V^{2}$ which prevents the two populations from mixing.
@@
In the present work we do not address the problem of obtaining uniform BV estimates for $u_{\varepsilon}$} and $v_{\varepsilon}$}, which appears to be a very challenging and worthwhile problem in its own right. Instead, we prove directly that the nonlinear terms $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$ and $v_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$ converge weakly to $u\,\partial_{x}p(s)$ and $v\,\partial_{x}p(s)$, respectively; that is, we identify their weak limits. We use the approach of compensated compactness developed by Tartar and Murat in the context of hyperbolic conservation laws [38, 47]. This is formulated in the language of Young measures (introduced in Section 4.1) which encode weak limits of nonlinear functions. We remark that the approach has been successfully applied (among many others) to hyperbolic conservation laws [22, 16], fast-reaction limits in reaction-diffusion systems [41, 46] or regularization of forward-backward diffusion problems [42, 46]. In contrast to these works, the Young measure considered here captures oscillations of two sequences $\{u_{\varepsilon}/s_{\varepsilon}\}$ and $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$ that do not converge strongly and there is no trivial relationship between them. Consequently, significant new developments are required in the way the compensated compactness framework is exploited. We present more detailed sketch of the proof in Section 2. Last but not least, we remark that Young measures are not only used to identify the limit, but also to define the concept of solutions (measure-valued solutions), where the limit cannot be identified [15, 8, 27].
@@
  • $\partial_{x}V^{i}$, $\partial_{x}K^{i,j}$, $\partial^{2}_{x}V^{i}$, $\partial^{2}_{x}K^{i,j}$ are in $L^{\infty}(\mathbb{R})$ and $\|\partial_{x}K^{i,j}\,\mathds{1}_{|x|\geq R}\|_{L^{1}(\mathbb{R})}\to 0$ as $R\to\infty$,
  • the initial conditions $u^{0},v^{0}\in L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ and $u^{0}\,|x|^{2},v^{0}\,|x|^{2}\in L^{1}(\mathbb{R})$ (for $\alpha>\frac{1}{3}$) or $u^{0}\,|x|,v^{0}\,|x|\in L^{1}(\mathbb{R})$ (for $\alpha\leq\frac{1}{3}$).
@@
We say that a pair $u,v\in L^{\infty}(0,T;L^{1}(\mathbb{R}))$ with $s:=u+v$ is a weak solution to (1.2) with initial conditions $u^{0},v^{0}\in L^{1}(\mathbb{R})$ if for all $\varphi,\phi\in C_{c}^{\infty}([0,T)\times\mathbb{R})$, $\partial_{x}s^{\alpha}\in L^{2}((0,T)\times\mathbb{R})$ and we have
@@
Under Assumption 1.1, there exists a weak solution to (1.2) in the sense of Definition 1.2 which satisfies $u,v\in L^{\infty}(0,T;L^{p}(\mathbb{R}))$ for all $p\in[1,\infty]$, $\partial_{x}s^{\frac{\alpha}{2}}\in L^{2}((0,T)\times\mathbb{R})$ and $s\,|x|^{2}\in L^{\infty}(0,T;L^{1}(\mathbb{R}))$ (if $\alpha>\frac{1}{3}$) or $s\,|x|\in L^{\infty}(0,T;L^{1}(\mathbb{R}))$ if $\alpha\leq\frac{1}{3}$. It is obtained as a limit of solutions to (1.3). More precisely, there exists a subsequence (not relabelled) such that
@@
  • $u_{\varepsilon}\to u$, $v_{\varepsilon}\to v$ strongly in $L^{p}(\mathcal{D})$ for all $p\in[1,\infty)$ and a.e. on $\mathcal{D}$, where the set $\mathcal{D}$ is defined as $\mathcal{D}:=\{(t,x)\in~[0,T]~\times~\mathbb{R}~:s(t,x)=0\mbox{ or }s(t,x)>0,\mathcal{V}^{1}[u,v]\neq\mathcal{V}^{2}[u,v]\}$,
  • $\partial_{x}s_{\varepsilon}^{{\alpha}}\to\partial_{x}s^{{\alpha}}$ strongly in $L^{2}((0,T)\times\mathbb{R})$.
The same result can be proved on a periodic domain $\mathbb{T}$.
@@
Structure of the paper. Section 2 contains a detailed sketch of the proof. Section 3 gathers the preparatory steps. We first present the standard a priori estimates for (1.3) in Section 3.1. In Section 3.2 we establish the corresponding weak compactness results. Sections 3.3 and 3.4 introduce two new conservation laws, while in Section 3.5 we quantify the defect measure $|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|^{2}$ as $\varepsilon\to 0$. Section 4 develops the more advanced tools required for the proof of Theorem 1.3. In Section 4.1 we introduce Young measures while in Section 4.2 we discuss compensated compactness, refining some known results. In Section 4.3 we reduce the proof of the existence of a weak solution to showing that a certain integral involving the Young measure vanishes. The proof of this final statement together with the proof of Theorem 1.3 is given in Section 5 and is divided into three steps, carried out in Sections 5.1, 5.2 and 5.3. For $\alpha\leq 2$, the second step can be proved by a simpler argument, presented in Section 5.4.
@@ 2. Sketch of the proof
$\partial_{t}U_{\varepsilon}=\partial_{x}F_{\varepsilon}+R_{\varepsilon},\qquad\partial_{t}V_{\varepsilon}=\partial_{x}G_{\varepsilon}+S_{\varepsilon},\qquad\mbox{ on }(0,T)\times\mathbb{R},$
@@
where $E_{\varepsilon}=u_{\varepsilon}\log u_{\varepsilon}+v_{\varepsilon}\log v_{\varepsilon}$, $h_{\varepsilon}\to 0$ in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ and $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}((0,T)\times\mathbb{R})$. We apply compensated compactness again as explained above (although now the remainder is not compact in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$; see below) and, after some careful algebra based mainly on the fact that $\lambda_{2}=\mathcal{F}\,\lambda_{1}$, we deduce that the Young measure is in fact a Dirac mass concentrated at $(0,0)$. When $\alpha\leq 2$, we can use another conservation law (see Section 3.4)
@@
As mentioned above, the compensated compactness tools used to exploit (2.2) or (2.3) need to be refined. The main difficulty is that the sequence $\{g_{\varepsilon}\}$ on the (RHS) is not compact in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, since $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ does not embed compactly into $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ in two dimensions. In the theory of hyperbolic conservation laws, a standard way to overcome this difficulty is to apply Murat’s lemma [25, Corollary 1, p.8], which relies on the fact that the terms on the left-hand sides of (2.2) and (2.3) are derivatives of bounded functions. However, this approach is not applicable here, since $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$ is bounded only in $L^{2}((0,T)\times\mathbb{R})$ and it seems difficult to improve its integrability. Instead, we weaken the assumptions in the compensated compactness framework to allow for remainder terms that are merely bounded in $\mathcal{M}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, exploiting the fact that the first equation enjoys better regularity. More precisely, in the PDE for $u_{\varepsilon}$, the remainder term has the form $\varepsilon\,\partial_{x}^{2}u_{\varepsilon}$, which is compact not only in $H^{-1}((0,T)\times\mathbb{R})$ but also in $L^{2}(0,T;H^{-1}(\mathbb{R}))$. Since the argument involves products of terms coming from different equations, the improved regularity in one equation compensates for the weaker regularity in the other. We refer to Theorem 4.3 for the precise statement and to Section 4.2 for a more detailed discussion of this issue.

### 16 Apr 2026 v2

@@ Abstract.
We consider cross-diffusion systems describing evolution of two species $u$ and $v$ moving according to Darcy’s law with the pressure law $p(s)=\frac{1}{\alpha-1}s^{\alpha-1}$ where $s=u+v$. One of the most challenging questions in the field is the construction of solutions to the problem in the presence of additional advection fields, without imposing any artificial structure on the fields or the initial conditions. Although advection arises naturally in these models, it breaks the symmetry of the system and prevents application of techniques developed in recent years. Here, we provide a new approach to construct solutions in one space dimension that works in a unified manner for all pressure exponents $\alpha\in(0,\infty)$ and for arbitrary initial data. In particular, in the regime $\alpha>1$, this yields the first existence result of its kind, obtained without any structural assumptions. We construct the solutions as a limit of a vanishing viscosity approximation $(u_{\varepsilon},v_{\varepsilon})$. The main challenge is to identify the limit of $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$, where $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$. The key new insight is that possible oscillations of $u_{\varepsilon}$ and $\partial_{x}p(s_{\varepsilon})$ are correlated, simplifying the Young measure analysis in the compensated compactness argument and allowing identification of the limit. Somewhat surprisingly, in contrast to the theory of $2\times 2$ hyperbolic systems, the argument relies on only three entropy-entropy flux pairs. This is particularly useful for $\alpha>2$, where it is unclear whether additional entropies are available.
@@ 1. Introduction
Cross-diffusion systems have attracted considerable attention in recent years due to their analytical challenges [25, 40, 20, 6, 35] and their applications in the applied sciences [16, 30, 44]. These partial differential equations, first formulated in the 1970s and 1980s [34, 8, 53, 9], still present many challenging and open mathematical problems. The aim of the present paper is to address one of them, namely, the existence of weak solutions to the following cross-diffusion system:
@@
describing the evolution of two species on $\mathbb{R}$ having nonnegative densities $u$, $v$, where
@@
$p(s)=\begin{cases}\frac{1}{\alpha-1}s^{\alpha-1}&\mbox{ if }\alpha\neq 1,\\ \log s&\mbox{ if }\alpha=1,\end{cases}\qquad\qquad\mbox{ for }\alpha\in(0,\infty),$ (1.2)
and it models the fact that species avoid overcrowded regions. The case $\alpha>1$ seems to be more relevant from the point of view of applications as the total density $s$ is expected to move with finite speed of propagation by an analogy to the porous medium equation [59, Chap. 14]. On the other hand, for $\alpha\leq 1$ one expects that $s$ is immediately globally supported as for heat equation or fast-diffusion equation, see [36, Theorem 2.4] and [58]. Moreover, the so-called incompressible limit $\alpha\to\infty$ shows that these models are related to the free boundary models of tissue growth for large $\alpha$ [47, 33].
@@
The target of this work is to construct weak solutions without imposing any structural conditions on $V^{i}$, $K^{i,j}$ or conditions guaranteeing that $u$, $v$ or $u+v$ are strictly positive. In particular, we do not make any assumption on the support of the initial conditions, the pressure law exponent and velocity fields (except mild regularity assumptions which could be probably improved). We also do not include the self-diffusion terms $\partial^{2}_{x}u$, $\partial^{2}_{x}v$ in (1.3).
@@
with $s_{\varepsilon}=u_{\varepsilon}+v_{\varepsilon}$ and study the limit $\varepsilon\to 0$. The difficult terms are nonlinear ones, namely $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$, $v_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$. Without advection (i.e. $\mathcal{V}^{1}=\mathcal{V}^{2}=0$) the system has attracted a large group of researchers and the understanding is now quite complete in arbitrary number of dimensions. In particular, there are several techniques to construct weak solutions: Aronson-Benilan estimates [4] can be used to obtain a uniform bound on $\partial^{2}_{x}p(s_{\varepsilon})$ [35], BV estimates on $\frac{u_{\varepsilon}}{s_{\varepsilon}}$, $\frac{v_{\varepsilon}}{s_{\varepsilon}}$ coming from a transport equation satisfied by these quantities lead to strong compactness of $u_{\varepsilon}$ and $v_{\varepsilon}$ in one dimension (probably this argument was used in [7] for the first time but see [11, 12] as well) while direct energy considerations show that $\{\partial_{x}p(s_{\varepsilon})\}$ is strongly compact in $L^{2}((0,T)\times\mathbb{R}^{d})$ for $\alpha\geq 1$ [23, 50, 28]. In contrast, for a single equation rather than a system, the problem with advection is well understood [24].
@@
  • [37] constructs weak solutions to (1.3) with $\alpha>1$ without nonlocal interaction $K^{i,j}=0$. The initial conditions are initially separated and the Authors impose an inequality between the velocities $\partial_{x}V^{1}$ and $\partial_{x}V^{2}$, which prevents the two populations from mixing.
@@
In the present work we do not address the problem of obtaining uniform BV estimates for $u_{\varepsilon}$} and $v_{\varepsilon}$}, which appears to be a very challenging and worthwhile problem in its own right. Instead, we prove directly that the nonlinear terms $u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$ and $v_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})$ converge weakly to $u\,\partial_{x}p(s)$ and $v\,\partial_{x}p(s)$, respectively; that is, we identify their weak limits. We use the approach of compensated compactness developed by Tartar and Murat in the context of hyperbolic conservation laws [45, 56]. This is formulated in the language of Young measures (introduced in Section 4.1) which encode weak limits of nonlinear functions. The general idea is to derive additional PDEs satisfied by nonlinear functions of the solution, the so-called entropy-entropy flux pairs. Applying the Div-Curl Lemma (see Section 4.2) then yields constraints on the Young measure which, provided sufficiently many such pairs are available, allow one to deduce convergence. We remark that the approach has been successfully applied (among many others) to hyperbolic conservation laws [26, 38, 39], mixed-type elliptic-hyperbolic systems [19], fast-reaction limits in reaction-diffusion systems [48, 54], and regularization of forward-backward diffusion problems [49, 54]. Last but not least, we remark that Young measures are not only used to identify the limit, but also to define the concept of solutions (measure-valued solutions), where the limit cannot be identified [18, 10, 31].
In contrast to the aforementioned works, the present paper requires significant new developments in how the compensated compactness framework is exploited. The Young measure considered here captures oscillations of two sequences, $u_{\varepsilon}/s_{\varepsilon}$} and $\partial_{x}s_{\varepsilon}^{\alpha}$}, which do not converge strongly and for which there is no simple relationship (for instance, via the divergence). Therefore, the system under consideration should be compared with $2\times 2$ hyperbolic conservation laws [26, 38, 39], where the argument typically relies on countably many entropy-entropy flux pairs. In our setting, such a family exists (Remark 3.9), but only for $\alpha\leq 2$ (Remark 3.8), which excludes many cases of interest; in particular, large values of $\alpha$ are important in tumor-growth models [47, 33]. Consequently, we carry out a careful analysis of the identities obtained via compensated compactness in order to complete the argument. A more detailed sketch of the proof is provided in Section 2.
@@
  • $\partial_{x}V^{i}$, $\partial_{x}K^{i,j}$, $\partial^{2}_{x}V^{i}$, $\partial^{2}_{x}K^{i,j}$ are in $L^{\infty}(\mathbb{R})$,
  • the initial conditions $u^{0},v^{0}\in L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ and $u^{0}\,|x|,v^{0}\,|x|\in L^{1}(\mathbb{R})$.
@@
We say that a pair $u,v\in L^{\infty}(0,T;L^{1}(\mathbb{R}))$ with $s:=u+v$ is a weak solution to (1.3) with initial conditions $u^{0},v^{0}\in L^{1}(\mathbb{R})$ if $\partial_{x}s^{\alpha}\in L^{2}((0,T)\times\mathbb{R})$ and for all $\varphi,\phi\in C_{c}^{\infty}([0,T)\times\mathbb{R})$ we have
@@
Under Assumption 1.1, there exists a weak solution to (1.3) in the sense of Definition 1.2 which satisfies $u,v\in L^{\infty}(0,T;L^{p}(\mathbb{R}))$ for all $p\in[1,\infty]$, $\partial_{x}s^{\frac{\alpha}{2}}\in L^{2}((0,T)\times\mathbb{R})$ and $s\,|x|\in L^{\infty}(0,T;L^{1}(\mathbb{R}))$. It is obtained as a limit of solutions to (1.4). More precisely, there exists a subsequence (not relabelled) such that
@@
  • $\partial_{x}s_{\varepsilon}^{{\alpha}}\to\partial_{x}s^{{\alpha}}$ strongly in $L^{2}((0,T)\times\mathbb{R})$,
  • $u_{\varepsilon}\to u$, $v_{\varepsilon}\to v$ strongly in $L^{p}(\mathcal{S})$ for all $p\in[1,\infty)$ and a.e. on $\mathcal{S}$, where the set $\mathcal{S}:=[0,T]~\times~\mathbb{R}\setminus\{(t,x):s(t,x)>0\mbox{ and }\partial_{x}\mathcal{V}^{1}[u,v]=\partial_{x}\mathcal{V}^{2}[u,v]\}$.
The same result can be proved on a periodic domain $\mathbb{T}$. It is also not difficult to extend this result to systems with additional growth terms of the form
$\begin{split}\partial_{t}u&=\partial_{x}(u\,\partial_{x}p(s))+\partial_{x}(u\,\partial_{x}\mathcal{V}^{1}[u,v])+u\,G^{1}(s),\\ \partial_{t}v&=\partial_{x}(v\,\partial_{x}p(s))+\partial_{x}(v\,\partial_{x}\mathcal{V}^{2}[u,v])+v\,G^{2}(s),\end{split}$ (1.6)
where $G^{i}:[0,\infty)\to\mathbb{R}$ and $G^{i}\in L^{\infty}(\mathbb{R})$. In tissue-growth processes, the additional source term acts to inhibit cell proliferation at high pressure levels [47]. We stress, however, that we cannot handle the case where $G$ depends on the individual species $u$ and $v$, as our method does not yield strong compactness of $\{u_{\varepsilon}\}$ and $\{v_{\varepsilon}\}$, unless the set $\mathcal{S}$ is the whole space $0,T]\times\mathbb{R}$ (this is the case, for instance, when $\partial_{x}\mathcal{V}^{1}=-1$ and $\partial_{x}\mathcal{V}^{2}=1$). We briefly outline the necessary modifications required to extend Theorem 1.3 to equation (1.6) in Section 5.5.
Remark 1.4.
We may also consider Darcy’s law (1.2) with the pressure depending not on the total density $s=u+v$, but instead on a weighted sum $\widetilde{s}=c_{u}\,u+c_{v}\,v$. This formulation is motivated by the fact that pressure may reflect physical characteristics such as cell size. For instance, five cells of type $u$ may occupy the same volume as ten cells of type $v$; see [15]. This case can be reduced to the standard setting $c_{u}=c_{v}=1$. Indeed, if $u,v$ solve the system where the pressure depends on the weighted sum $\widetilde{s}$ with kernels $\widetilde{K^{i,j}}$ and velocities $\widetilde{V^{i}}$, then $c_{u}\,u,c_{v}\,v$ solve the original system (1.1) with kernels ${K^{i,1}}=\frac{1}{c_{u}}\widetilde{K^{i,1}},{K^{i,2}}=\frac{1}{c_{v}}\widetilde{K^{i,2}}$ (the velocity fields $\widetilde{V^{i}}$ do not change).
@@
Structure of the paper. Section 2 contains a detailed sketch of the proof. Section 3 gathers the preparatory steps. We first present the standard a priori estimates for (1.4) in Section 3.1. In Section 3.2 we establish the corresponding weak compactness results. Sections 3.3 and 3.4 introduce two new conservation laws, while in Section 3.5 we quantify the defect measure $|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|^{2}$ as $\varepsilon\to 0$. Section 4 develops the more advanced tools required for the proof of Theorem 1.3. In Section 4.1 we introduce Young measures while in Section 4.2 we discuss the compensated compactness approach. In Section 4.3 we reduce the proof of the existence of a weak solution to showing that a certain integral involving the Young measure vanishes. The proof of this final statement together with the proof of Theorem 1.3 is given in Section 5 and is divided into three steps, carried out in Sections 5.1, 5.2 and 5.3. For $\alpha\leq 2$, the second step can be proved by a simpler argument, presented in Section 5.4. Finally, in Section 5.5, we outline the necessary modifications required to extend the reasoning to the nonconservative case (1.6).
@@ 2. Sketch of the proof
$\partial_{t}U_{\varepsilon}=\partial_{x}F_{\varepsilon}+R_{\varepsilon},\qquad\partial_{t}V_{\varepsilon}=\partial_{x}G_{\varepsilon}+S_{\varepsilon}\qquad\mbox{ on }(0,T)\times\mathbb{R},$
@@
where $E_{\varepsilon}=u_{\varepsilon}\log u_{\varepsilon}+v_{\varepsilon}\log v_{\varepsilon}$, $h_{\varepsilon}\to 0$ in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ and $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}((0,T)\times\mathbb{R})$. We apply compensated compactness again as explained above (although now the remainder is not compact in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ and one has to use Theorem 4.3). The resulting identity is technically challenging to analyze due to its nonlinear structure arising from the logarithmic terms. We therefore perform several decompositions and apply delicate estimates to the identity obtained via compensated compactness. Combining these with the relation $\lambda_{2}=\mathcal{F}\,\lambda_{1}$, we conclude that the Young measure is in fact a Dirac mass concentrated at $(0,0)$. When $\alpha\leq 2$, we can use another conservation law (see Section 3.4)